UT 20131025 - 20131027 LUCI2 Daytime Commissioning

Observer: KPolsterer, APramskiy, PBuschkamp, Wseifert (remote)
Telescope Operator:
Software Support:
LBTO Support: JPower, ESolheid

Summary

Issue with Rack Cooling Control solved

Scripts to take flexure data work

CalUnit service still needs to be debugged, ie CalUnit must at present be moved in manually

Plan for the period

Solve issue with Rack Cooling Control by Alexander, Jenny, Michael & Walter
Tests of calibration unit service, tests of scripts to take flexure data by Alexander & Kai
Further tests of MOS by Peter

Preparations

Details

25 Oct 2013:

Jennifer swapped the Rack Coltrol Unit => everything is working fine now.
Alex tested the flexure scripts which involve movement of the rotator => works fine now!
Alex did also some tests with the CalUnit Service. Apparently there are still some bugs to be fixed.

26 Oct 2013:

-- JochenHeidt - 26 Oct 2013
Topic revision: r3 - 26 Oct 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback